The door frames used by abuilder of one standard height 1.830 m the hights of men are normally distributed with mean 1.730 m ansfadard deviation 0.064 m. a) find the proportion of men that will be taller the door frame .b) find the frame height such that one man in a thousand will be taller then the frame height.
1
Expert's answer
2020-06-08T20:28:44-0400

a) "P(X>1.830)=P(Z>\frac{1.830-1.730}{0.064}=P(Z>1.56)=0.0594."


b) "P(Z>z)=0.001"

"P(Z<z)=0.999"

"z=3.09"

"\frac{x-1.730}{0.064}=3.09"

"x=3.09*0.064+1.730=1.928\; m."